OpenStudy (anonymous):

simplify the logarithm log2 (64)

OpenStudy (anonymous):

To what power must 2 be taken to in order to get 64? 2^0 = 1 2^1 = 2 2^2 = 4 2^3 = 8 2^4 = 16 ... @clarkandrea1964

OpenStudy (anonymous):

2^6 gives you 64

OpenStudy (anonymous):

\[\log2_{2^6} \]

OpenStudy (anonymous):

1/6\log2_2

OpenStudy (anonymous):

1/6 should be your answer

OpenStudy (anonymous):

I think Log2 (64) is 6 but I am not sure \[\log_{2} (64) is 6 but I am \not sure \]

thomaster (thomaster):

lol just 2^6 xD

OpenStudy (anonymous):

oh I see, thank you!

OpenStudy (anonymous):

LOL--I should have known that